OpenStudy (anonymous):

Limit as X approaches 4 [sin(6x)]/x

OpenStudy (amistre64):

sin(u) = u - u^3/3! + u^5/5! - u^7/7! + ... sin(6x) = 6x - (6x)^3/3! + (6x)^5/5! - (6x)^7/7! + ... sin(6x)/x = 6 - 6^2 x/3! + 6^5 x^4/5! - 6^7 x^6/7! + ... hmmm

OpenStudy (amistre64):

seems to me to be simpler to just plug in x=4 since there is no divide by 0 from it

OpenStudy (anonymous):

Ia m sorry, i wrote it wrong, it is x approaching 0

OpenStudy (anonymous):

My bad.

OpenStudy (amistre64):

oh, then the power series is useful since its a polynomial; all the x terms zero out and we are left with the first term

OpenStudy (amistre64):

does [ ] indicate something like a step function, or is that just your way of seperating the top from the bottom?

OpenStudy (anonymous):

its just my way to separate it

OpenStudy (amistre64):

k, then since: sin(6x)/x = 6 - 6^2 x/3! + 6^5 x^4/5! - 6^7 x^6/7! + ... when x=0 we have sin(6(0))/0 = 6 - 0 + 0 - 0 + 0 - 0 + 0 - .... = 6

OpenStudy (anonymous):

since it is all divided by zero, wouldn't the answer be zero?

OpenStudy (anonymous):